How mdraw fits in with TEX

     software 'other'                           preview dvi2vdu printer .ln3 dvi2ln3                               .tex mdr2cgm                                         TeX vregis mdraw                                             .dvi .pic          .tex .mdr .cgm

Fig. 1

If you are a person who uses TEX, and you would like to include some simple graphics along with you document, or would like to see what the document looks like before you print it out, today is your lucky day.

The drawing above shows the relationship of several programs, the ones in the polygons, and the types of files they use and produce. Normally TEX produces a ‘.dvi’ file which could only be sent to the printer. Now there is a new program to convert the ‘.dvi’ for display on a terminal. There is also new input into TEX (actually into dvi2ln3) in the form of a ‘.pic’ file. This file is basically a screen dump of graphics drawn by mdraw. In order to easily use this ‘.pic’ file in a TEX file, another TEX file is created which holds all the bizarre commands to handle the ‘.pic’ file. All that needs to be done to the top document is to include this supporting TEX file.